1200*28=1200*4+n*32 solve calculations how to calculate
 

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
1200*28 = 1200*4 + n*32
1200*28 -1200*4 = n*32
1200(28-4) = n*32
1200*24 = n*32
n = 1200*24/32
n = 1200*3/4
n = 300*3
n = 900
 
Regards
Arun (askIITians forum expert)